Find moment of inertia of solid cylinder through an axis which passes through middle of its width
1
Expert's answer
2020-10-21T06:32:58-0400

Explanations & Calculations


  • Consider radius = r, height = h, mass = M for the cylinder.
  • Moment of inertia of about its central axis is "\small \bold{\frac{1}{2}Mr^2}"

Finding that about an axis at "\large\frac{h}{2}" : passes through middle of width

  • For that consider an elemental disc of mass "\small \delta m"
  • The easiest way of deriving the final solution is to consider the moment of inertia of it through the axis in consideration.
  • For that one should know MI of a solid plate through its middle perpendicular axis as "\frac{1}{2}\delta m r^2"
  • And by the law for MI of perpendicular axes, MI of that plate through an axis lie at any diameter as "\small \frac{1}{4}\delta m r^2"
  • Finally MI of that plate about an axis lies a distance of x parallel to that previously considered diameter as "\small \frac{1}{4}\delta m r^2 +\delta mx^2"

Then considering this elemental MI of that plate about the axis considered for the whole cylinder and integrating over the full height, answer could be obtained.

"\qquad\qquad
\begin{aligned}
\small \delta m &= \small \pi r^2\delta x\times \rho \,\,\,\,\,\,\,\,\,: \rho\text{ = volumetric mass}\\
&= \small \pi r^2\delta x\times \frac{M}{\pi r^2h}\\
&= \small \frac{M}{h}\delta x\\
\\
\small I &= \small \int_{-h/2}^{h/2}\delta I\\
&= \small \small \int_{-h/2}^{h/2} \frac{1}{4}\Big(\frac{M}{H}\delta x\Big)r^2+\Big(\frac{M}{h}\delta x\Big) x^2 \\
&= \small \frac{Mr^2}{4h} \int_{-h/2}^{h/2} \delta x + \frac{M}{h} \int_{-h/2}^{h/2} x^2 \delta x\\
& = \small \frac{Mr^2}{4h} [x]_{-h/2}^{h/2} + \frac{M}{h} [\frac{x^3}{3}]_{-h/2}^{h/2}\\
&= \small \frac{Mr^2}{4h} [h] + \frac{M}{h} [\frac{h^3}{12}]\\
\small I &= \small \bold{M\Big[\frac{r^2}{4}+\frac{h^2}{12}\Big]}
\end{aligned}"